# LightRetriever: A LLM-based Text Retrieval Architecture with Extremely Faster Query Inference

### v4

@@ A.4 Additional ablation about the effect of auxiliary KL loss
Previous works for BERT-based asymmetric retrievers [74, 33] emphasize the importance of alignment between asymmetric students and full symmetric teachers. To explore the effect of such alignment, besides contrastive loss as the main loss function, we also explore to adopt the KL loss as an auxiliary to align the asymmetric similarity scores $S^{student}=v_{q}^{den}\cdot v_{d}^{den}$ to full-sized symmetric similarity scores $S^{teacher}=Enc_{q}(q)\cdot Enc_{d}(d)$,
@@
A.5 Training datasets
Dataset infos
Following the settings of previous LLM-based retrievers [72, 5, 48], our work uses 23 training sets pre-built by tDRO [48] or Sentence Transformers Training Data [60], including 20 English and 3 Chinese datasets. Detailed dataset information is listed in Table 6 as follows.
@@ A.5 Training datasets
Amazon Review (2018) [54] English Amazon $999$ 0.01 0.58%
cMedQA2 [82] Chinese Chinese Medical 99936 1 5.84%
DuReader [58] Chinese Chinese Web Collections 86366 1 5.05%
Eli5 [20] English Reddit 325390 0.05 0.95%
Fever [67] English Wikipedia QA 109808 0.8 5.13%
FiQA [49] English Financial 5498 1 0.32%
GooQA Pairs [31] English Web Collections 3012347 0.025 4.40%
HotpotQA [81] English Wikipedia QA 85000 0.5 2.48%
MSMARCO [53] English Web Collections 502854 1 29.39%
NFCorpus [7] English Medical 2585 1 0.15%
NPR [45] English News 594376 0.05 1.74%
NQ [37] English Wikipedia QA 58800 1 3.44%
PQA Pairs [41] English Wikipedia QA 99999 0.5 2.92%
Quora Duplicates Triples [26] English Forum Duplicates 97011 0.5 2.83%
S2ORC Title-Abstract [43] English Semantic Scholar 99998 0.5 2.92%
SciFact [70] English S2ORC 806 1 0.05%
SPECTER [13] English Semantic Scholar 136642 0.25 2.00%
StackExchange Duplicates (Title-Body) [17] English Forum Duplicates 250516 0.25 3.66%
E5 Synthetic [4] English GPT Synthetic Data 224791 1 13.14%
T2Ranking [78] Chinese Chinese Web Collections 200362 0.5 5.86%
Trivia [29] English Wikipedia QA 60370 0.5 1.76%
Wikihow [34] English WikiHow 128543 0.25 1.88%
Yahoo Answers (Title-Answer) [83] English Yahoo 1198018 0.05 3.50%
Many downstream retrieval benchmarks [66, 77] are evaluated out-of-domain, where no training samples from the downstream sources are available. To avoid the overfitting of training sets, as is shown in the above Table 6, we slightly tune the epoch used in the dataset sampling, by following previous works in [55, 72, 48].
Training set decontamination
Training set decontamination is essential for developing fair, unbiased, and robust retrieval systems. Existing training datasets are derived from multiple sources, whose training data have possible duplicates with the test sets. Following tDRO [48], we perform strict training set decontamination with SimHash [50] to ensure no training queries appear in the BeIR [66] and CMTEB Retrieval [77]. The original sizes, deduplicated sizes, and duplicate numbers as listed in the Table 7.
@@
A.6 Task instructions
As previously mentioned in Section 2.1, modern LLM-based retrievers [72, 5] introduce task-specific instructions to format input queries, which enables higher retrieval capacities. These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation [56]. Following Mistral-E5 [72], we reuse most of its task instructions, and format the instruction and query in the following format:
@@ A.6 Task instructions
The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral [72] and tDRO [48] for reproducibility. Note that one training set may correspond to multiple instructions to avoid overfitting.
@@
    GooQA Pairs: Given a web search query, retrieve the corresponding answers from Google
@@
A.7 Efficiency optimizations techniques
@@
The customized causal mask ensures that the token blocks can attend to the common prompt area, but not attend to each other. Let the sequence length be $L$, prompt length be $P$, micro block (each token + <eos>) width be $w$, and let $q\in[0,L-1]$ denote the query index, $k\in[0,L-1]$ the key index. The attention mask (0 means attend, -$\infty$ means masked) is defined as,
@@
, where the vocabulary dimension $V$ is much larger than the hidden dimension $H$. For example, for Llama3.1-8b [68] model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$. This means that, given batch size B, sequence length T, hidden size H, vocabulary size V, the intermediate tensors with shape $[B,T,V]$ in the projection, ReLU, and Log are 31x larger than a dense tensor $[B,T,H]$, which consumes huge amounts of GPU memory during training and inference. However, only a tensor with shape $[B,V]$ is needed for the final aggregated representations. Thus, an optimization for computing the sparse vector is needed urgently.
@@
Proper optimizations are important. For example, we found it hard to reproduce quickly with some of the open-sourced baselines [5], which is roughly 6x slower than our implementation when testing with the same size of LLM. Our experiments involve training with 23 train sets and inference with another 23 test sets, which requires proper optimizations urgently. We import other optimizations based on the implementations  [^3] Please refer to [48]’s implementation on https://github.com/ma787639046/tdro. released by tDRO [48]. These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC [42]; 2) Sequence packing [35] with Flash Attention 2 [18], which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel [24], which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache [23] for optionally enlarging batch sizes if necassary.
A.8 Detailed main results

### v5

@@ A.5 Additional ablation about the effect of auxiliary KL loss
Previous works for BERT-based asymmetric retrievers (Wang and Lyu, 2023; Kim et al., 2024) emphasize the importance of alignment between asymmetric students and full symmetric teachers. To explore the effect of such alignment, besides contrastive loss as the main loss function, we also explore to adopt the KL loss as an auxiliary to align the asymmetric similarity scores $S^{student}=v_{q}^{den}\cdot v_{d}^{den}$ to full-sized symmetric similarity scores $S^{teacher}=Enc_{q}(q)\cdot Enc_{d}(d)$,
@@
Amazon Review (2018) (Ni et al., 2019) English Amazon $999$ 0.01 0.58%
cMedQA2 (Zhang et al., 2018) Chinese Chinese Medical 99936 1 5.84%
DuReader (Qiu et al., 2022) Chinese Chinese Web Collections 86366 1 5.05%
Eli5 (Fan et al., 2019) English Reddit 325390 0.05 0.95%
Fever (Thorne et al., 2018) English Wikipedia QA 109808 0.8 5.13%
FiQA (Maia et al., 2018) English Financial 5498 1 0.32%
GooAQ Pairs (Khashabi et al., 2021) English Web Collections 3012347 0.025 4.40%
HotpotQA (Yang et al., 2018b) English Wikipedia QA 85000 0.5 2.48%
MSMARCO (Nguyen et al., 2016) English Web Collections 502854 1 29.39%
NFCorpus (Boteva et al., 2016) English Medical 2585 1 0.15%
NPR (Lucas et al., 2023) English News 594376 0.05 1.74%
NQ (Kwiatkowski et al., 2019) English Wikipedia QA 58800 1 3.44%
PQA Pairs (Lewis et al., 2021) English Wikipedia QA 99999 0.5 2.92%
Quora Duplicates Triples (Iyer et al., 2017) English Forum Duplicates 97011 0.5 2.83%
S2ORC Title-Abstract (Lo et al., 2020) English Semantic Scholar 99998 0.5 2.92%
SciFact (Wadden et al., 2020) English S2ORC 806 1 0.05%
SPECTER (Cohan et al., 2020) English Semantic Scholar 136642 0.25 2.00%
StackExchange Dups (Title-Body) (da Silva et al., 2018) English Forum Duplicates 250516 0.25 3.66%
E5 Synthetic (BeastyZ, 2024) English GPT Synthetic Data 224791 1 13.14%
T2Ranking (Xie et al., 2023) Chinese Chinese Web Collections 200362 0.5 5.86%
Trivia (Joshi et al., 2017) English Wikipedia QA 60370 0.5 1.76%
Wikihow (Koupaee and Wang, 2018) English WikiHow 128543 0.25 1.88%
Yahoo Answers (Title-Answer) (Zhang et al., 2015) English Yahoo 1198018 0.05 3.50%
A.6 Training datasets
Dataset infos
Following the settings of previous LLM-based retrievers (Wang et al., 2024a; BehnamGhader et al., 2024; Ma et al., 2025), our work uses 23 training sets pre-built by tDRO (Ma et al., 2025) or Sentence Transformers Training Data, including 20 English and 3 Chinese datasets. Detailed dataset information is listed in Table 11 as follows.
Many downstream retrieval benchmarks (Thakur et al., 2021; Xiao et al., 2024) are evaluated out-of-domain, where no training samples from the downstream sources are available. To avoid the overfitting of training sets, as is shown in the above Table 11, we slightly tune the epoch used in the dataset sampling, by following previous works in (Ni et al., 2022; Wang et al., 2024a; Ma et al., 2025).
@@ A.6 Training datasets
Training set decontamination
Training set decontamination is essential for developing fair, unbiased, and robust retrieval systems. Existing training datasets are derived from multiple sources, whose training data have possible duplicates with the test sets. Following tDRO (Ma et al., 2025), we perform strict training set decontamination with SimHash (Manku et al., 2007) to ensure no training queries appear in the BeIR (Thakur et al., 2021) and CMTEB Retrieval (Xiao et al., 2024). The original sizes, deduplicated sizes, and duplicate numbers as listed in the Table 12.
A.7 Task instructions
As previously mentioned in Section 2.1, modern LLM-based retrievers (Wang et al., 2024a; BehnamGhader et al., 2024) introduce task-specific instructions to format input queries, which enables higher retrieval capacities. These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation (Ouyang et al., 2022). Following Mistral-E5 (Wang et al., 2024a), we reuse most of its task instructions, and format the instruction and query in the following format:
@@ A.7 Task instructions
The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral (Wang et al., 2024a) and tDRO (Ma et al., 2025) for reproducibility. Note that one training set may correspond to multiple instructions to avoid overfitting.
@@
    GooAQ Pairs: Given a web search query, retrieve the corresponding answers from Google
@@
A.8 Efficiency optimizations techniques
@@
The customized causal mask ensures that the token blocks can attend to the common prompt area, but not attend to each other. Let the sequence length be $L$, prompt length be $P$, micro block (each token + eos) width be $w$, and let $q\in[0,L-1]$ denote the query index, $k\in[0,L-1]$ the key index. The attention mask (0 means attend, -$\infty$ means masked) is defined as,
@@
, where the vocabulary dimension $V$ is much larger than the hidden dimension $H$. For example, for Llama3.1-8b (Touvron et al., 2023) model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$. This means that, given batch size B, sequence length T, hidden size H, vocabulary size V, the intermediate tensors with shape $[B,T,V]$ in the projection, ReLU, and Log are 31x larger than a dense tensor $[B,T,H]$, which consumes huge amounts of GPU memory during training and inference. However, only a tensor with shape $[B,V]$ is needed for the final aggregated representations. Thus, an optimization for computing the sparse vector is needed urgently.
@@
Proper optimizations are important. For example, we found it hard to reproduce quickly with some of the open-sourced baselines (BehnamGhader et al., 2024), which is roughly 6x slower than our implementation when testing with the same size of LLM. Our experiments involve training with 23 train sets and inference with another 23 test sets, which requires proper optimizations urgently. We import other optimizations based on the implementations released by tDRO (Ma et al., 2025). These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC (Li et al., 2023); 2) Sequence packing (Kundu et al., 2024) with Flash Attention 2 (Dao, 2024), which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel (Hsu et al., 2024), which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache (Gao et al., 2021) for optionally enlarging batch sizes if necassary.
A.9 Detailed main results
@@
A.10 Additional results on CoIR and BRIGHT
Benchmarks of our query-lightweight design on CoIR (Code Retrieval) and BRIGHT (Reasoning-intensive Retrieval) have been conducted using three backbones (Llama-3.2-1B, Llama-3.1-8B, Qwen-2.5-7B). Results are shown below in Table 17 and 18.
As code retrieval and reasoning-intensive retrieval are more challenging out-of-domain (OOD) tasks requiring deep query or structure understanding, our proposed query-lightweight encoders are less performant compared to strong LLM embedding models like E5-Mistral (-6.2pp on CoIR, -6.3pp on BRIGHT), which is expected in our work. Our method is more suitable for normal relevance-based retrieval. Queries requiring complicated instructions or code understanding are not suitable to be applied to the query-lightweight encoder.
Table 17: Retrieval performance (nDCG@10) on the CoIR benchmark.
Model Apps CosQA SyntheticText2SQL CodeSearchNet CodeSNCC CodeTransContest CodeTransDL StackOverFlowQA CodeFeedBackMT CodeFeedBackST Avg
BM25 1.0 14.0 16.9 26.8 34.7 50.1 8.7 56.8 34.7 54.3 29.8
BGE-M3 7.4 22.7 48.8 43.2 47.6 47.9 31.2 61.0 33.5 49.9 39.3
E5-Mistral 21.3 31.3 66.0 54.3 65.3 82.6 33.2 91.5 33.7 72.7 55.2
LightRetriever
Llama3.2-1B 11.2 30.1 52.0 60.1 58.6 57.7 27.9 77.8 29.6 64.5 47.0
Llama3.1-8B 18.1 29.5 54.5 57.8 59.0 59.7 25.7 82.0 33.8 68.1 48.8
Qwen2.5-7B 18.4 30.0 54.0 57.7 61.8 61.5 22.3 84.4 33.1 66.3 49.0
Table 18: Retrieval performance (nDCG@10) on the BRIGHT benchmark.
Model Bio. Earth. Econ. Psy. Rob. Stack. Sus. Leet. Pony AoPS TheT. TheQ. Avg
BM25 18.9 27.2 14.9 12.5 13.6 18.4 15.0 24.4 7.9 6.2 10.4 4.9 14.5
BGE-large 11.7 24.6 16.6 17.5 11.7 10.8 13.3 26.7 5.7 6.0 13.0 6.9 13.7
E5-Mistral 18.6 26.0 15.5 15.8 16.3 11.2 18.1 28.7 4.9 7.1 26.1 26.8 17.9
LightRetriever
Llama3.2-1B 13.7 21.4 11.2 11.1 18.5 13.0 12.7 15.2 4.0 2.9 6.5 6.0 11.4
Llama3.1-8B 14.2 22.3 12.6 11.1 17.1 13.9 15.0 16.6 8.1 0.7 4.0 5.7 11.8
Qwen2.5-7B 19.3 24.0 13.8 13.1 14.1 11.3 14.3 11.8 5.2 1.1 5.8 5.7 11.6